OpenStudy (anonymous):

In the porblem, t is a real number and P = (x, y) is the point on the unit circle that corresponds to t. Find the exact value of the indicated trigonometric function of t.

OpenStudy (anonymous):

1)\[\left(\frac{ 4 }{ 7 } \frac{ \sqrt{33} }{ 7 } \right) \] find sin t.